It is estimated that 30% of all drivers have some kind of medical aid in South Africa. What is the probability that in a sample of 10 drivers:
1.1 At least 2 will have a medical aid
1.2. More than 9 will have a medical aid
1
Expert's answer
2019-04-18T02:56:03-0400

The binomial probability formula is the result of getting exactly k successes in n number of trials.


"P(k \ success\ in\ n \ trials)=\begin{pmatrix}
 n \\
 k
\end{pmatrix}p^k(1-p)^{n-k}"

We have that


"n=10, p=0.3"

1.1 At least 2 will have a medical aid 


"P(X\geq 2)=1-(P(X=0)+P(X=1))="

"=1-\begin{pmatrix}
 10 \\
 0
\end{pmatrix}0.3^0(1-0.3)^{10-0}-\begin{pmatrix}
 10 \\
 1
\end{pmatrix}0.3^1(1-0.3)^{10-1}\approx 0.850692"

1.2. More than 9 will have a medical aid


"P(X> 9)=P(X=10)=\begin{pmatrix}
 10 \\
 10
\end{pmatrix}0.3^{10}(1-0.3)^{10-10}\approx 0.000006"